Aman’s age is three times his son’s age. Ten years ago he was five times his son’s age. Find their present ages. 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let the age of Aman's son be x then age of Aman will be 3x.
According to the question
5(x-10)=3x-10
⇒5x-50=3x-10
⇒5x-3x=-10+50
⇒2x=40
⇒x=20
Aman's son age = x = 20 years Aman age = 3x = 3×20 = 60 years

Thanks